OpenStudy (anonymous):

Multiplying a number by which of the following would result in a smaller number? A 10^-1 B 10^0 C 10^1 D 10^2

OpenStudy (lgbasallote):

Hint!!! A) 10^-1 = 1/10 B) 10^0 = 1 C) 10^1 = 10 D) 10^2 = 100

OpenStudy (lgbasallote):

so which one is the smallest @samyflashy ?

OpenStudy (anonymous):

A, THANKS :)